# Logarithmic Regret for Online KL-Regularized Reinforcement Learning

### v2

@@ 1 Introduction
However, despite the empirical success, a fundamental connection between KL regularization and learning efficiency has not been established so far theoretically. While some recent studies have attempted to analyze these frameworks, they often rely on standard analysis techniques, yielding regret guarantees similar to those of standard RL (Xiong et al., 2024a; Xie et al., 2024; Xiong et al., 2024b; Cen et al., 2024), or they depend on strong coverage assumptions and restricted to the bandit settings (Zhao et al., 2024). A notable exception is the approach of Tiapkin et al. (2023, 2024), which achieves $\widetilde{\mathcal{O}}(H^{5}d^{2}\eta/\epsilon)$ sample complexity for KL-regularized linear MDPs, and $\widetilde{\mathcal{O}}(H^{5}S^{2}A\eta/\epsilon)$ sample complexity for tabular MDPs. However, their analysis focuses on the pure exploration or best policy identification setting, where the goal is to find near-optimal policies using the least possible amount of interactions with the environment, rather than the online setting where the agent needs to trade off between exploration and exploitation. Therefore, The following pivotal question remains open:
@@ Appendix A Proofs for KL-Regularized Contextual Bandits
$\leq\eta\mathbb{E}_{x\sim d_{0}}\bigl{[}\sum_{a\in\mathcal{A}}\bigl{(}\widehat{R}_{t-1}(x,a)+b_{t-1}(x,a)-R^{*}(x,a)\bigr{)}\cdot\underbrace{\bigl{(}f_{\lambda}(x,a)-R^{*}(x,a)\bigr{)}\cdot\pi_{f_{\lambda}}^{\eta}(a|x)}_{U(x,a)}\bigr{]},$ (A.3)
@@
To bound the function $U(f,x,a)$ on the right-hand side of (A.3), we take the derivative as follows:
$\frac{\partial\bigl{(}f(x,a)-R^{*}(x,a)\bigr{)}\cdot\pi_{f}^{\eta}(a|x)}{\partial f(x,a)}$
$=\pi_{f}^{\eta}(a|x)+\eta\bigl{(}f(x,a)-R^{*}(x,a)\bigr{)}\cdot\Big{(}\pi_{f}^{\eta}(a|x)-\pi_{f}^{\eta}(a|x)^{2}\Big{)}\geq 0,$
where we use (A.5) in the first equality and use $\pi_{f}^{\eta}(a|x)\leq 1$ in the inequality. Therefore, since $f_{\lambda}(x,a)\leq\widehat{R}_{t-1}(x,a)+b_{t-1}(x,a)$ for any $(x,a)\in\mathcal{X}\times\mathcal{A}$, we further have the following bound for RHS of (A.3):

### v3

@@ 1 Introduction
However, despite the empirical success, a fundamental connection between KL regularization and learning efficiency has not been established so far theoretically. While some recent studies have attempted to analyze these frameworks, they often rely on standard analysis techniques, yielding regret guarantees similar to those of standard RL (Xiong et al., 2024a; Xie et al., 2024; Xiong et al., 2024b; Cen et al., 2024), or they depend on strong coverage assumptions and restricted to the bandit settings (Zhao et al., 2024). The following pivotal question remains open:
@@ Appendix A Proofs for KL-Regularized Contextual Bandits
$\leq\eta\mathbb{E}_{x\sim d_{0}}\bigl{[}\underbrace{\sum_{a\in\mathcal{A}}\bigl{(}\widehat{R}_{t-1}(x,a)+b_{t-1}(x,a)-R^{*}(x,a)\bigr{)}\cdot\bigl{(}f_{\lambda}(x,a)-R^{*}(x,a)\bigr{)}\cdot\pi_{f_{\lambda}}^{\eta}(a|x)}_{U(\lambda)}\bigr{]},$ (A.3)
@@
We use the notation $\delta_{t}=\widehat{R}_{t}+b_{t}-R^{*}$. Let
$U(\lambda)=\sum_{a\in\mathcal{A}}\lambda\delta_{t-1}(x,a)^{2}\cdot\frac{\pi_{0}(a|x)\cdot\exp\big{(}\eta(\lambda\delta_{t-1}(x,a)+R^{*}(x,a))\big{)}}{\mathbb{E}_{\pi_{0}}\exp\big{(}\eta(\lambda\delta_{t-1}(x,a)+R^{*}(x,a))\big{)}},\quad\text{for}~{}\lambda\in[0,1].$
We take the derivative as follows:
$\frac{\partial U(\lambda)}{\partial\lambda}=$ $\sum_{a\in\mathcal{A}}\Big{[}\delta_{t-1}(x,a)^{2}\pi_{f_{\lambda}}^{\eta}+\lambda\delta_{t-1}(x,a)^{2}\cdot\Big{(}\frac{\pi_{0}(a|x)\exp\big{(}\eta(\lambda\delta_{t-1}(x,a)+R^{*}(x,a))\big{)}\cdot\eta\delta_{t-1}(x,a)}{\mathbb{E}_{\pi_{0}}\exp\big{(}\eta(\lambda\delta_{t-1}(x,a)+R^{*}(x,a))\big{)}}\Big{)}$
$\qquad-\frac{\pi_{0}(a|x)\exp\big{(}\eta(\lambda\delta_{t-1}(x,a)+R^{*}(x,a))\big{)}\cdot\mathbb{E}_{\pi_{0}}\eta\delta_{t-1}(x,a)\cdot\exp\big{(}\eta(\lambda\delta_{t-1}(x,a)+R^{*}(x,a))\big{)}}{\big{(}\mathbb{E}_{\pi_{0}}\exp\big{(}\eta(\lambda\delta_{t-1}(x,a)+R^{*}(x,a))\big{)}\big{)}^{2}}\Big{]}$
$=$ $\sum_{a\in\mathcal{A}}\Big{[}\delta_{t-1}(x,a)^{2}\pi_{f_{\lambda}}^{\eta}+\lambda\delta_{t-1}(x,a)^{2}\pi_{f_{\lambda}}^{\eta}\cdot\eta\big{(}\delta_{t-1}(x,a)-\mathbb{E}_{\pi_{f_{\lambda}}^{\eta}}\delta_{t-1}(x,a)\big{)}\Big{]}$
$=$ $\mathbb{E}_{\pi_{f_{\lambda}}^{\eta}}\big{[}\delta_{t-1}(x,a)^{2}\big{]}+\lambda\Big{\{}\mathbb{E}_{\pi_{f_{\lambda}}^{\eta}}\big{[}\delta_{t-1}(x,a)^{3}\big{]}-\mathbb{E}_{\pi_{f_{\lambda}}^{\eta}}\big{[}\delta_{t-1}(x,a)^{2}\big{]}\cdot\mathbb{E}_{\pi_{f_{\lambda}}^{\eta}}\big{[}\delta_{t-1}(x,a)\big{]}\Big{\}}$
$\geq$ $0,$
where the inequality holds since for random variable $X\geq 0$ we have
$\mathbb{E}[X^{3}]-\mathbb{E}[X^{2}]\cdot\mathbb{E}[X]$ $=\mathbb{E}[(X^{2}-\mathbb{E}[X^{2}])(X-\mathbb{E}[X])]$
$=\mathbb{E}[\big{(}X^{2}-(\mathbb{E}[X])^{2}\big{)}(X-\mathbb{E}[X])]+\mathbb{E}[\big{(}(\mathbb{E}[X])^{2}-\mathbb{E}[X^{2}]\big{)}(X-\mathbb{E}[X])]$
$=\mathbb{E}[\big{(}X+\mathbb{E}[X]\big{)}(X-\mathbb{E}[X])^{2}]+0$
$\geq 0.$
Therefore, since $f_{\lambda}(x,a)\leq\widehat{R}_{t-1}(x,a)+b_{t-1}(x,a)$ for any $(x,a)\in\mathcal{X}\times\mathcal{A}$, we can obtain that
$U(\lambda)\leq U(1)=\sum_{a\in\mathcal{A}}\delta_{t-1}(x,a)^{2}\cdot\pi_{\widehat{R}_{t-1}+b_{t-1}}^{\eta}(a|x).$
We further have the following bound for RHS of (A.3):